OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 1,024 and a4 = -16?

OpenStudy (campbell_st):

well 1st the formula for a term in a geometric series is \[a_{n} = a_{1} \times r^{n - 1}\] 1st you need to find the common ratio before finding the 7th term so use what you know \[-16 = 1024 \times r^{4 -1}\] solve for r. when you have it, go to the formula and find the 7th term with n = 7

OpenStudy (anonymous):

r=-1/4?

OpenStudy (campbell_st):

thats correct... so just evaluate \[a_{7} = 1024 \times (\frac{-1}{4})^{7 -1}\]

OpenStudy (campbell_st):

you should get a positive answer

OpenStudy (anonymous):

so it is .25

OpenStudy (campbell_st):

thats right

OpenStudy (anonymous):

Thankyou that helped alot